# Learning to Deblur Polarized Images

### v1

@@ Abstract
A polarization camera can capture four polarized images with different polarizer angles in a single shot, which is useful in polarization-based vision applications since the degree of polarization (DoP) and the angle of polarization (AoP) can be directly computed from the captured polarized images. However, since the on-chip micro-polarizers block part of the light so that the sensor often requires a longer exposure time, the captured polarized images are prone to motion blur caused by camera shakes, leading to noticeable degradation in the computed DoP and AoP. Deblurring methods for conventional images often show degenerated performance when handling the polarized images since they only focus on deblurring without considering the polarization constrains. In this paper, we propose a polarized image deblurring pipeline to solve the problem in a polarization-aware manner by adopting a divide-and-conquer strategy to explicitly decompose the problem into two less ill-posed sub-problems, and design a two-stage neural network to handle the two sub-problems respectively. Experimental results show that our method achieves state-of-the-art performance on both synthetic and real-world images, and can improve the performance of polarization-based vision applications such as image dehazing and reflection removal.
@@ 1 Introduction
Polarization is an important property of light in addition to its amplitude and phase. By virtue of the physical information encoded in the polarization-relevant parameters (e.g., the degree of polarization (DoP) and the angle of polarization (AoP) of the incoming light to the sensor), it has been widely introduced into the field of computer vision (Lyu et al, 2019, 2022; Zhou et al, 2021; Mei et al, 2022; Deschaintre et al, 2021), aiming to conquer the bottleneck of conventional vision algorithms. With the development of the division of focal plane (DoFP) technology, modern polarization cameras (e.g., Lucid Vision Phoenix polarization camera [^1] https://thinklucid.com/product/phoenix-5-0-mp-polarized-model/) can capture four spatially-aligned and temporally-synchronized polarized images with different polarizer angles in a snapshot, bringing convenience to the acquisition of the DoP and AoP. However, the captured polarized images are prone to motion blur caused by camera shakes, since the on-chip micro-polarizers block part of the light so that the sensor often requires a longer exposure time. Therefore, it is of practical interest to deblur polarized images for acquiring the DoP and AoP accurately.
@@
    Polarization constraints have to be maintained across multiple images in addition to removing the motion blur, making the problem highly ill-posed. Deblurring methods for conventional images can only process the polarized images in a polarization-unaware manner.
@@
    The DoFP technology adopted by modern polarization cameras induces higher level of sensor noise (compared to the same sensor without micro-polarizers) (Kumar et al, 2007; Qiu et al, 2019; Abubakar et al, 2018), leading to degradation of low-level features (e.g., structural and boundary information).
@@
    Since the on-chip micro-polarizers attenuate scene radiance and the extent of attenuation is scene-dependent (Wu et al, 2020; Zhou et al, 2023a), the exposures turn to be spatially-variant and arbitrary, making high-level features (e.g., semantic and contextual information) less distinctive.
@@
    Blur resistance property of Stokes parameters: two (out of three) Stokes parameters have similar appearance to the gradient distribution of the corresponding unpolarized image and are less affected by motion blur.
@@
Based on these inspiring characteristics of polarized images, we specially build a two-stage neural network to handle the two sub-problems respectively: The first stage is an unpolarized image estimator that leverages the less-contaminated structural and boundary information encoded in the Stokes parameters $\mathbf{S}_{1,2}$ for repairing degraded low-level features. The second stage is a polarized image reconstructor that transfers the semantic and contextual information from the estimated unpolarized image to the polarized images with robustly extracted high-level features. To summarize, this paper makes contributions by demonstrating three customized model designs to deal with the three issues mentioned above:
@@
    leveraging the less-contaminated structural and boundary information encoded in the Stokes parameters $\mathbf{S}_{1,2}$ for repairing degraded low-level features; and
@@ 2 Related work
2.1 Deblurring unpolarized images.
@@
Generally, the methods that can deblur unpolarized images could be divided into two categories: image deblurring methods and video deblurring methods. We will have an overview on them respectively in the following.
@@
Image deblurring methods. Previous works turned the problem into a maximum a posteriori (MAP) estimation problem and adopted numerical optimization to enforce sharpness (Fergus et al, 2006; Xu et al, 2013; Pan et al, 2016; Yang and Ji, 2019; Chen et al, 2020, 2021a; Zhang et al, 2022c). Recently, learning-based approaches have been introduced to deal with this problem (Zhang et al, 2022b). By extracting features from a large amount of training data to reduce the ill-posedness, they often show better performance and higher efficiency than numerical optimization-based methods. Some works proposed to restore the sharp image from its blurry counterpart in an end-to-end manner using neural networks with different kinds of architectures, including convolutional neural network (CNN) (Nah et al, 2017; Zhang et al, 2018a; Tao et al, 2018; Gao et al, 2019; Zhang et al, 2019; Suin et al, 2020; Cho et al, 2021; Chi et al, 2021; Mou et al, 2022; Ji et al, 2022; Chen et al, 2022), generative adversarial network (GAN) (Kupyn et al, 2018, 2019; Zhang et al, 2020; Wen et al, 2021b; Li et al, 2022a), multi-layer perceptron (MLP) (Tu et al, 2022), diffusion probabilistic model (DPM) (Whang et al, 2022), vision transformer (ViT) (Zamir et al, 2022; Wang et al, 2022b; Tsai et al, 2022; Chu et al, 2022), etc., demonstrating visually impressive results. Some works proposed to deblur in a two-stage manner, by adopting neural networks to estimate blur-related physical parameters (e.g., blur kernels and motion flows) first and then utilizing them to guide the deblurring process (Kaufman and Fattal, 2020; Ren et al, 2020; Dong et al, 2021; Tran et al, 2021; Chen et al, 2021b; Chakrabarti, 2016; Sun et al, 2015; Chen et al, 2018; Gong et al, 2017; Yuan et al, 2020), showing good generalization ability. However, they are not suitable for handling multiple polarized images since they can only deblur in a polarization-unaware manner by processing one polarized image at a time, without considering the polarization constraints.
@@
Video deblurring methods. Video deblurring methods can process multiple frames instead of a single frame at a time. Early works adopted numerical optimization to perform blur kernel estimation and deconvolution for restoring sharp frames (Bar et al, 2007; Wulff and Black, 2014; Ren et al, 2017). With the development of deep-learning, neural networks have also been adopted to handle this problem. These methods could be divided into three categories: window-based methods, recurrent methods, and memory-based methods. Window-based methods take multiple frames as input and try to reconstruct the middle frame, by referencing neighbouring frames within a small fixed-size window (Su et al, 2017; Pan et al, 2020; Suin and Rajagopalan, 2021; Zhang et al, 2022a). Recurrent methods deblur frames in sequence, by collectively aggregating frame features till up-to-date to provide restoration cues (Wieschollek et al, 2017; Hyun Kim et al, 2017; Nah et al, 2019; Zhou et al, 2019; Zhong et al, 2020; Li et al, 2021; Zhu et al, 2022; Jiang et al, 2022; Wang et al, 2022a). Memory-based methods allow the feature propagation from the entire sequence when restoring the given frame, by using memory banks to provide appropriate spatio-temporal information for each region and exploiting the features of all frames (Ji and Yao, 2022). However, they are still not suitable for handling multiple polarized images since the polarized images do not contain temporal information as the video frames.
2.2 Enhancing polarized images.
@@
There are many methods aiming to enhance the polarized images (e.g., joint chromatic and polarimetric demosaicing (Tyo et al, 2009; Qiu et al, 2019; Wen et al, 2019, 2021a; Pistellato et al, 2022; Liu et al, 2022), denoising (Zhang et al, 2017; Abubakar et al, 2018; Tibbs et al, 2018; Li et al, 2020; Liang et al, 2022; Li et al, 2023), and low-light enhancement (Hu et al, 2020; Li et al, 2022b; Xu et al, 2022; Zhou et al, 2023b)) to acquire the DoP and AoP accurately for improving the performance of downstream applications of polarization-based vision. Although these methods are specially designed for processing multiple polarized images, the enhancement tasks targeted by them have completely different properties from the deblurring problem.
@@
Figure 1: Architecture of the neural network that implements the proposed polarized image deblurring pipeline. It consists of two stages: an unpolarized image estimator that uses $\hat{\mathbf{S}}_{1,2}$ as priors to estimate the sharp unpolarized image $\mathbf{I}_{\text{guide}}$, and a polarized image reconstructor that reconstructs $\mathbf{I}_{\alpha_{1,2,3,4}}$ under the guidance of $\mathbf{I}_{\text{guide}}$. All images are normalized for visualization.
@@ 3 Polarized image deblurring pipeline
Denoting the total intensity of the incoming light to the sensor as $\mathbf{I}$ (i.e., the unpolarized image captured without polarizers), when placing a polarizer in front of the camera, the captured polarized image $\mathbf{I}_{\alpha}$ can be calculated using Malus’ law (Hecht et al, 2002):
@@
where $\alpha$ is the polarizer angle, $\mathbf{p}\in[0,1]$ and $\bm{\theta}\in[0,\pi]$ denote the DoP and AoP of the incoming light to the sensor respectively. Note that in this paper, we assume the camera response function to be linear since polarization cameras usually output images with a linear camera response function, and focus on linear polarization (i.e., do not consider circular polarization) since polarization cameras only equip linear polarizers.
By virtue of four-directional, on-chip micro-polarizers provided by the DoFP technology, a modern polarization camera can capture four spatially-aligned and temporally-synchronized polarized images $\mathbf{I}_{\alpha_{1,2,3,4}}$ with different polarizer angles $\alpha_{1,2,3,4}=0^{\circ},45^{\circ},90^{\circ},135^{\circ}$ in a snapshot, which brings convenience to the acquisition of $\mathbf{p}$ and $\bm{\theta}$. This is because $\mathbf{p}$ can be calculated using
@@
$\mathbf{B}_{\alpha_{i}}=\mathbf{I}_{\alpha_{i}}\otimes\mathbf{K}+\mathbf{N}(i=1,2,3,4),$ (6)
where $\otimes$ stands for the convolution operation, $\mathbf{K}$ denotes a spatially-variant blur kernel determined by the motion field, and $\mathbf{N}$ represents a noise term depending on the input signal and camera configuration. Note that both $\mathbf{B}$ and $\mathbf{B}_{\alpha_{i}}$ ($i=1,2,3,4$) can be considered to share the same blur kernel $\mathbf{K}$, since the pixel shifts among the polarized images caused by mosaicing is small enough (usually 1 pixel for a $4\times 4$ mosaicing of polarization and color filters inside a polarization camera), which will not cause significant changes to the motion field and is often ignored by other works designed for enhancing the polarized images (Hu et al, 2020; Li et al, 2022b; Xu et al, 2022; Zhou et al, 2023b).
We aim to restore $\mathbf{p}$ and $\bm{\theta}$ from their degenerated counterparts $\hat{\mathbf{p}}$ and $\hat{\bm{\theta}}$ by deblurring $\mathbf{B}_{\alpha_{1,2,3,4}}$. Unlike conventional image deblurring problems which only need to focus on removing the blurry artifacts, our problem becomes more ill-posed since we need to consider how to maintain the polarization constraints across the polarized images in addition to deblurring. This is because according to Equation (2), Equation (3), and Equation (4), the relationship among $\mathbf{p}$, $\bm{\theta}$, and $\mathbf{I}_{\alpha_{1,2,3,4}}$ involves non-linear operations, which makes the accuracy of $\mathbf{p}$ and $\bm{\theta}$ highly sensitive to the maintenance of the polarization constraints (Hu et al, 2020; Zhou et al, 2023b). Therefore, we adopt a divide-and-conquer strategy to explicitly decompose the problem into two less ill-posed sub-problems, and propose a two-stage polarized image deblurring pipeline to handle them respectively in a polarization-aware manner:
@@
Figure 2: Left: Two of the Stokes parameters ($\mathbf{S}_{1,2}$) have similar “appearance” to the gradient distribution $\mathbf{G}$ of the corresponding unpolarized image (please zoom-in for better details). Middle: LPIPS (Zhang et al, 2018b) scores between the blurry/sharp variable pairs, including $\mathbf{S}_{1,2}$ and $\mathbf{G}$. Right: Semantic distance between $\mathbf{I}_{\alpha_{1,2,3,4}}$ and $\mathbf{I}$, $\mathbf{I}_{\alpha_{1,2,3,4}}$ and $\mathbf{B}_{\alpha_{1,2,3,4}}$

### v2

@@ Abstract
A polarization camera can capture four linear polarized images with different polarizer angles in a single shot, which is useful in polarization-based vision applications since the degree of linear polarization (DoLP) and the angle of linear polarization (AoLP) can be directly computed from the captured polarized images. However, since the on-chip micro-polarizers block part of the light so that the sensor often requires a longer exposure time, the captured polarized images are prone to motion blur caused by camera shakes, leading to noticeable degradation in the computed DoLP and AoLP. Deblurring methods for conventional images often show degraded performance when handling the polarized images since they only focus on deblurring without considering the polarization constraints. In this paper, we propose a polarized image deblurring pipeline to solve the problem in a polarization-aware manner by adopting a divide-and-conquer strategy to explicitly decompose the problem into two less ill-posed sub-problems, and design a two-stage neural network to handle the two sub-problems respectively. Experimental results show that our method achieves state-of-the-art performance on both synthetic and real-world images, and can improve the performance of polarization-based vision applications such as image dehazing and reflection removal.
@@ 1 Introduction
Polarization is an important property of light in addition to its amplitude and phase. By virtue of the physical information encoded in the polarization-relevant parameters (e.g., the degree of linear polarization (DoLP) and the angle of linear polarization (AoLP) of the incoming light to the sensor), it has demonstrated unique advantages in various computer vision tasks, such as glass segmentation (Mei et al, 2022), color constancy (Ono et al, 2022), and shape reconstruction (Deschaintre et al, 2021). Besides, it also plays a crucial role in applications for autonomous driving, such as road detection (Li et al, 2020a), reflection removal (Lyu et al, 2019, 2022), dehazing (Zhou et al, 2021), and HDR reconstruction (Zhou et al, 2023a), conquering the bottleneck of conventional vision algorithms in this field.
With the development of the division of focal plane (DoFP) technology, modern polarization cameras (e.g., Lucid Vision Phoenix polarization camera [^1] https://thinklucid.com/product/phoenix-5-0-mp-polarized-model/) can capture four spatially-aligned and temporally-synchronized linear polarized images with different polarizer angles in a snapshot, bringing convenience to the acquisition of the DoLP and AoLP. However, the captured polarized images are prone to motion blur caused by camera shakes, since the on-chip micro-polarizers block part of the light so that the sensor often requires a longer exposure time. Therefore, it is of practical interest to deblur polarized images for acquiring the DoLP and AoLP accurately.
@@
    Polarization constraints (the numerical relationship between multiple polarized images constrained by the physical meanings of polarization) have to be preserved across multiple images in addition to removing the motion blur, making the problem highly ill-posed. Since the deblurring methods designed for conventional images can only process the polarized images in a polarization-unaware manner (i.e., treat multiple polarized images as ordinary frames and ignore the physical meanings of polarization), they have no chance to take the preservation of polarization constraints into consideration.
@@
    The DoFP technology adopted by modern polarization cameras induces a higher level of sensor noise (compared to the same sensor without micro-polarizers) (Kumar et al, 2007; Qiu et al, 2019; Abubakar et al, 2018), leading to degradation of low-level features (e.g., structural and boundary information). Since current deblurring methods usually aim to deal with the images captured by conventional sensors, they cannot take advantage of the information and priors provided by the polarization-related parameters to mitigate the difficulties during the feature extraction and fusion procedure.
@@
    Since the on-chip micro-polarizers attenuate scene radiance and the extent of attenuation is scene-dependent (Wu et al, 2020; Zhou et al, 2023a), the exposures turn out to be spatially-variant and arbitrary, making high-level features (e.g., semantic and contextual information) less distinctive. Since recent deblurring methods usually adopt a single-stage pipeline that depends strongly on the quality of high-level features to hallucinate the sharp contents to restore the sharp image from its blurry counterpart in an end-to-end manner, they tend to suffer from ringing artifacts and fake textures.
@@
    Blur resistance property of Stokes parameters: two (out of three) Stokes parameters (which are polarization-related parameters and will be introduced later) have similar appearance to the gradient distribution of the corresponding unpolarized image and are less affected by motion blur.
@@
Based on these inspiring characteristics of polarized images, we specially build a two-stage neural network to handle the two sub-problems respectively: The first stage is an unpolarized image estimator that leverages the less-contaminated structural and boundary information encoded in the Stokes parameters for repairing degraded low-level features. The second stage is a polarized image reconstructor that transfers the semantic and contextual information from the estimated unpolarized image to the polarized images with robustly extracted high-level features. To summarize, this paper makes contributions by demonstrating three customized model designs to deal with the three issues mentioned above:
@@
    leveraging the less-contaminated structural and boundary information encoded in the Stokes parameters for repairing degraded low-level features; and
@@ 2 Related work
2.1 Deblurring unpolarized images
@@
Generally, methods for deblurring unpolarized images can be divided into two categories: image deblurring and video deblurring methods. We will have an overview on them respectively in the following.
@@
Image deblurring methods. Previous works turned the problem into a maximum a posteriori (MAP) estimation problem and adopted numerical optimization to enforce sharpness (Fergus et al, 2006; Xu et al, 2013; Pan et al, 2016; Yang and Ji, 2019; Chen et al, 2020, 2021a; Zhang et al, 2022c). Recently, learning-based approaches have been introduced to deal with this problem (Zhang et al, 2022b). By extracting features from a large amount of training data to reduce the ill-posedness, they often show better performance and higher efficiency than numerical optimization-based methods. Some works proposed to restore the sharp image from its blurry counterpart in an end-to-end manner using neural networks with different kinds of architectures, including convolutional neural network (CNN) (Nah et al, 2017; Zhang et al, 2018a; Tao et al, 2018; Gao et al, 2019; Zhang et al, 2019; Suin et al, 2020; Cho et al, 2021; Chi et al, 2021; Mou et al, 2022; Ji et al, 2022; Chen et al, 2022; Liu et al, 2024), generative adversarial network (GAN) (Kupyn et al, 2018, 2019; Zhang et al, 2020b; Wen et al, 2021b; Li et al, 2022a), multi-layer perceptron (MLP) (Tu et al, 2022), diffusion probabilistic model (DPM) (Whang et al, 2022; Ren et al, 2023), vision transformer (ViT) (Zamir et al, 2022; Wang et al, 2022b; Tsai et al, 2022; Chu et al, 2022; Kong et al, 2023), etc., demonstrating visually impressive results. Some works proposed to deblur in a two-stage manner, by adopting neural networks to estimate blur-related physical parameters (e.g., blur kernels and motion flows) first and then utilizing them to guide the deblurring process (Kaufman and Fattal, 2020; Ren et al, 2020; Dong et al, 2021; Tran et al, 2021; Chen et al, 2021b; Chakrabarti, 2016; Sun et al, 2015; Chen et al, 2018; Gong et al, 2017; Yuan et al, 2020), showing good generalization ability. Some methods adopted unrolling algorithms to bridge the gap between classical MAP optimization and pure deep learning techniques, which are also useful in deblurring or other image enhancement tasks (Kruse et al, 2017; Eboli et al, 2020; Zhang et al, 2020a). However, they are not suitable for handling multiple polarized images since they can only deblur in a polarization-unaware manner by processing one polarized image at a time, without considering the polarization constraints.
@@
Video deblurring methods. Video deblurring methods can process multiple frames instead of a single frame at a time. Early works adopted numerical optimization to perform blur kernel estimation and deconvolution for restoring sharp frames (Bar et al, 2007; Wulff and Black, 2014; Ren et al, 2017). With the development of deep-learning, neural networks have also been adopted to handle this problem. These methods could be divided into three categories: window-based methods, recurrent methods, and memory-based methods. Window-based methods take multiple frames as input and try to reconstruct the middle frame, by referencing neighboring frames within a small fixed-size window (Su et al, 2017; Pan et al, 2020; Suin and Rajagopalan, 2021; Zhang et al, 2022a). Recurrent methods deblur frames in sequence, by collectively aggregating frame features till up-to-date to provide restoration cues (Wieschollek et al, 2017a; Hyun Kim et al, 2017; Nah et al, 2019; Zhou et al, 2019; Zhong et al, 2020; Li et al, 2021; Zhu et al, 2022; Jiang et al, 2022; Wang et al, 2022a). Memory-based methods allow the feature propagation from the entire sequence when restoring the given frame, by using memory banks to provide appropriate spatio-temporal information for each region and exploiting the features of all frames (Ji and Yao, 2022). Additionally, there are methods designed for deblurring burst images (Aittala and Durand, 2018; Wieschollek et al, 2017b), which can also process multiple frames. However, they are still not suitable for handling multiple polarized images since the polarized images do not contain temporal information as the video frames or burst images.
2.2 Polarization and image enhancement
Here, we will introduce the intersection of polarization and image enhancement in two aspects: enhancing polarized images and polarization-based image enhancement.
@@ 2.2 Polarization and image enhancement
Enhancing polarized images. There are many methods aiming to enhance the polarized images (e.g., joint chromatic and polarimetric demosaicing (Tyo et al, 2009; Qiu et al, 2019; Wen et al, 2019, 2021a; Pistellato et al, 2022; Liu et al, 2022; Yi et al, 2024; Guo et al, 2024), denoising (Zhang et al, 2017; Abubakar et al, 2018; Tibbs et al, 2018; Li et al, 2020b; Liang et al, 2022; Li et al, 2023), and low-light enhancement (Hu et al, 2020; Li et al, 2022b; Xu et al, 2022; Zhou et al, 2023b)) to acquire the DoLP and AoLP accurately for improving the performance of downstream applications of polarization-based vision. Some methods perform enhancement in a direct manner, by processing multiple polarized images simultaneously in the intensity domain (e.g., IPLNet (Hu et al, 2020)). However, their performance may not that good, since enhancing in the intensity domain cannot make effective use of the information provided by polarization. On the contrary, some methods adopt a Stokes-domain enhancement pipeline (e.g., ColorPolarNet (Xu et al, 2022) and PLIE (Zhou et al, 2023b)) to enhance the polarized images in an indirect manner. These methods usually show higher performance and robustness, since the polarization characteristics of the incoming light to the sensor are fully encoded in the Stokes parameters, making the enhancement process more polarization-aware. Although the above methods are specially designed for processing multiple polarized images, the enhancement tasks targeted by them have completely different properties from the deblurring problem.
Polarization-based image enhancement. Different from enhancing polarized images, polarization-based image enhancement aims to utilize the polarization information to handle the image enhancement problems. Since the reflection and scattering phenomenons would alter the polarization property of light, polarization information can be used as priors to eliminate the reflection artifacts (Lyu et al, 2019; Lei et al, 2020; Lyu et al, 2022) or remove the haze (Zhou et al, 2021; Ma et al, 2024a, b) inside the images. Since the polarization cameras can not only offer spatially-variant exposures but also contain abundant structural and contextual information of the scene (encoded in the DoLP and AoLP), we can also use polarization information to achieve HDR reconstruction (Wu et al, 2020; Ting et al, 2021; Zhou et al, 2023a)). Although the above methods may also utilize the polarization-related parameters to provide guidance, their underlying strategies and objectives diverge from those of methods designed for enhancing polarized images, as they are designed to solve different types of problems.
@@
Figure 1: Architecture of the neural network that implements the proposed polarized image deblurring pipeline. It consists of two stages: an unpolarized image estimator that uses $\mathbf{S}^{*}_{1,2}$ (the blurry counterparts of the Stokes parameters $\mathbf{S}_{1,2}$) as priors to estimate the sharp unpolarized image $\mathbf{I}_{\text{guide}}$, and a polarized image reconstructor that reconstructs $\mathbf{I}_{\alpha_{1,2,3,4}}$ under the guidance of $\mathbf{I}_{\text{guide}}$. All images are normalized for visualization.
@@ 3 Polarized image deblurring pipeline
Denoting the total intensity of the incoming light to the sensor as $\mathbf{I}$ (i.e., the unpolarized image captured without polarizers), when placing a linear polarizer in front of the camera, the captured polarized image $\mathbf{I}_{\alpha}$ can be calculated using Malus’ law (Hecht et al, 2002):
@@
where $\alpha$ is the polarizer angle, $\mathbf{p}\in[0,1]$ and $\bm{\theta}\in[0,\pi]$ denote the DoLP and AoLP of the incoming light to the sensor respectively. Note that in this paper, we assume the camera response function to be linear since polarization cameras usually output images with a linear camera response function, and focus on linear polarization (i.e., do not consider circular polarization) since polarization cameras only equip linear polarizers.
By virtue of four-directional, on-chip micro-polarizers provided by the DoFP technology, a modern polarization camera can capture four spatially-aligned and temporally-synchronized linear polarized images $\mathbf{I}_{\alpha_{1,2,3,4}}$ with different polarizer angles $\alpha_{1,2,3,4}=0^{\circ},45^{\circ},90^{\circ},135^{\circ}$ in a snapshot, which brings convenience to the acquisition of $\mathbf{p}$ and $\bm{\theta}$. This is because $\mathbf{p}$ can be calculated using
@@
$\mathbf{B}_{\alpha_{i}}=\mathbf{I}_{\alpha_{i}}\otimes\mathbf{K}+\mathbf{N}_{i}(i=1,2,3,4),$ (6)
where $\otimes$ stands for the convolution operation, $\mathbf{K}$ denotes a spatially-variant blur kernel determined by the motion field, $\mathbf{N}$ and $\mathbf{N}_{i}$ represent noise terms depending on the input signal and camera configuration (Hu et al, 2020; Xu et al, 2022; Zhou et al, 2023b). Note that both $\mathbf{B}$ and $\mathbf{B}_{\alpha_{i}}$ ($i=1,2,3,4$) can be considered to share the same blur kernel $\mathbf{K}$, since the pixel shifts among the polarized images caused by mosaicing is small enough (usually 1 pixel for a $4\times 4$ mosaicing of polarization and color filters inside a polarization camera), which will not cause significant changes to the motion field and is often ignored by other works designed for enhancing the polarized images (Hu et al, 2020; Li et al, 2022b; Xu et al, 2022; Zhou et al, 2023b).
We aim to restore $\mathbf{p}$ and $\bm{\theta}$ from their degraded counterparts $\mathbf{p}^{*}$ and $\bm{\theta}^{*}$ by deblurring $\mathbf{B}_{\alpha_{1,2,3,4}}$. Unlike conventional image deblurring problems which only need to focus on removing the blurry artifacts, our problem becomes more ill-posed since we need to consider how to preserve the polarization constraints across the polarized images in addition to deblurring. This is because according to Equation (2), Equation (3), and Equation (4), the relationship among $\mathbf{p}$, $\bm{\theta}$, and $\mathbf{I}_{\alpha_{1,2,3,4}}$ involves non-linear operations, which makes the accuracy of $\mathbf{p}$ and $\bm{\theta}$ highly sensitive to the preservation of the polarization constraints (Hu et al, 2020; Zhou et al, 2023b). Therefore, we adopt a divide-and-conquer strategy to explicitly decompose the problem into two less ill-posed sub-problems, and propose a two-stage polarized image deblurring pipeline to handle them respectively in a polarization-aware manner: